1.what is the z-score of 40 in a math test when the mean is 58 with a standard deviation of 9?

1
Expert's answer
2022-03-25T06:58:43-0400

Z-score =( x- μ)/σ

where x is the random value =40

μ is the mean =58

σ is the standard deviation =9

Z= (40-58)/9 =-18/9 =-2

Z-score is -2